# Frequency Is What You Need: Word-frequency Masking Benefits Vision-Language Model Pre-training

### v2

@@ Abstract
Vision Language Models (VLMs) can be trained more efficiently if training sets can be reduced in size. Recent work has shown the benefits of masking text during VLM training using a variety of approaches: truncation, random masking, block masking and syntax masking. In this paper, we show that the best masking strategy changes over training epochs and that, given sufficient training epochs. We analyze existing text masking approaches including syntax masking, which is currently the state of the art, and identify the word frequency distribution as important in determining their success. Experiments on a large range of data sets demonstrate that syntax masking is outperformed by other approaches, given sufficient epochs, and that our proposed frequency-based approach, called Contrastive Language-Image Pre-training with Word Frequency Masking (CLIPF) has numerous advantages. The benefits are particularly evident as the number of input tokens decreases.
Keywords Vision-Language Model  $\cdot$ Multimodal Data  $\cdot$ Word Frequency Masking.
@@ 1 Introduction
However, pre-training CLIP models typically requires thousands of GPU-days [3, 9]. To accelerate the pre-training process, FLIP, A-CLIP, and CLIPA use masking strategies for image and text tokens when pre-training CLIP models [8, 11, 12]. In addition to the speed-up, masking makes it possible use large batch sizes, which enhances the benefits of contrastive learning and has been considered as a form of regularization, reducing the risk of overfitting [8]. For these reasons, despite the loss of some information from the original image and text, masking strategies have the potential to improve model performance.
@@
Figure 1: Zero-shot classification accuracy on ImageNet-1K over training epochs for CLIPF and CLIPA strategies. The backbone of the image encoder is ViT-B/16. The models are pre-trained on CC12M for 45 epochs with 75% text masking ratio. We also use image masking (75%) to speed up pre-training. All approaches are subsequently fine-tuned with one epoch on the full, unmasked data.
Currently, the state of the art in text masking is represented by CLIPA [12], which compares four masking strategies: truncation, random word selection, random block selection, and syntax-based selection, i.e., masking on the basis of part-of-speech (POS) information. The conclusion is that syntax masking is the strategy that, overall, provides the best quality. In this paper, we carry out a set of experiments that demonstrate that this conclusion must be revisited. Specifically, as shown in Fig. 1, if enough epochs are used, other forms of masking substantially outperform syntax masking. The conclusion of CLIPA [12] misses the weakness of syntax masking because it trains with a limited number of epochs (6.4 epochs).
The special focus of our paper is a frequency masking approach that we introduce: CLIPF (CLIP Frequency masking). In Fig. 1, we see that with enough epochs CLIPF and CLIPA block masking perform similarly. We provide an analysis that demonstrates why CLIPF, block masking, and random masking emerge as clearly superior to syntax masking and truncation. The analysis studies the distribution of words by frequency of occurrence in the training data after masking. It shows that good masking strategies maintain helpful properties of this distribution and that syntax masking distorts the distribution of words over POS categories. These observation leads to the main insight of our paper, namely, when masking text for VLM pre-training, what you need for best results is to pay attention to the impact of masking on word frequency.
Our CLIPF approach is useful in practice because it performs relatively well in the range of 16-40 epochs. This range is a typical number of epochs to train a system in practice. For example, Open_CLIP [9] trains on 30 epochs. In Fig. 1, we see that in this range CLIPF outperforms other masking strategies. We carry out a set of experiments with 30 epochs to shed further light on the role of frequency in masking by comparing CLIPF vs. the different variants of CLIPA masking on zero-shot tasks and on image retrieval. We also investigate training on a larger data set.
The history of text masking in language model training dates to [13], where subsampling of frequent words in the training text was used to improve the performance of skip-gram language models. The study of text masking for VLMs got off to a slow start, since initially [8] reported that text masking, although it promotes efficiency, deteriorated performance. The CLIPA [12] work is important because it expanded the types of masking studied and demonstrated that masking can indeed improve VLM performance. To our knowledge, no previous work has attempted to discover the reason that certain forms of text masking outperform others or identified the importance of the way in which text masking approaches impact word frequency.
In sum, our paper makes the following contributions to the study of text masking in pre-training VLMs:
@@
  • We show that word frequency distribution after masking differentiates between successful and less successful text masking strategies.
  • We carry out an in-depth comparison between existing masking strategies (CLIPA variants) and an approach that carries out masking by using frequency information directly, CLIPF. Our experiments show that CLIPF consistently outperforms syntax masking, previously considered the best choice for masking and provide a systematic view of how text masking can accelerate pre-training of VLMs while at the same time maintaining or improving VLM performance.
Our work is reproducible because it uses training data sets that are openly available and small enough to allow researchers with limited computational resources to repeat our experiments, and because we release our code [^1] https://github.com/MingliangLiang3/CLIPF.git.
@@ 2.1 Vision-Language Models
Vision-language models (VLMs), such as CLIP, ALIGN, LiT, and FLIP [3, 4, 10, 8], learn image embeddings from large-scale natural language supervision using contrastive learning [14], which is important for making models transferable. Because CLIP does not release its pre-training dataset, OpenCLIP pre-trained models on several large-scale public datasets [15, 16, 17, 18] to reproduce CLIP. This implementation adheres to the scaling law observed in Neural Language Models [19] and Vision-Language Models [3], where loss decreases according to a power-law relationship with model size, dataset size, and computational resources utilized during training. While large language models (LLMs) have advanced rapidly, we focus on vision-language models (VLMs) like CLIP, as they are essential tools for bridging visual and textual modalities in current research. Although CLIP and OpenCLIP are pre-trained on very large-scale datasets, recent works like DeCLIP [20] and SLIP [5] demonstrate that Vision-language Models (VLMs) can be trained efficiently with smaller datasets, enabling researchers with limited computational resources to contribute. Moreover, masking image patches or reducing image resolution during VLM pre-training also substantially decreases computational costs while maintaining or enhancing performance, as shown by techniques like FLIP, RECLIP, CLIPA [8, 21, 12].
@@ 2.2 Text masking for VLM training
Masking text presents another viable option to speed up the pre-training of VLMs, and is particularly effective alongside strategies that employ a high ratio of image patch masking. FLIP further explored text masking techniques, namely, random and truncation masking, to enhance efficiency [8].
While random text masking involves masking text arbitrarily, truncation masking in FLIP strategically eliminates padding tokens and retrains the words that preceded them.
@@
Subsequently, CLIPA [12] introduced block masking and syntax masking. As already mentioned in Section 1, CLIPA [12] shows that text masking during training can improve the performance of a VLM. CLIPA syntax masking prioritizes the retention of nouns, followed by adjectives and then other words. In [12], syntax masking outperforms the truncation, random, and block approaches. However, [12] does not vary the number of training epochs. An important contrast with our work, as previously mentioned, is our attention to the number of training epochs. Further, CLIPF, as we will see, counters the danger of giving overly high priority to nouns, leading to overfitting, and also avoids overly suppressing words of the POS category “Other”.
An initial effort to exploit frequency information for text masking during VLM training was made by SW-CLIP [22]. Unfortunately, SW-CLIPF introduces a frequency cutoff threshold, which reduces performance. In order to exploit frequency information effectively each training text needs to fill as many available input tokens as possible, as is done in CLIPF.
3 Text Masking
In this section, we describe the text masking strategies that we study in this paper: the current state of the art, namely, the text masking strategies of CLIPA [12] (Section 3.1) and our frequency masking strategy CLIPF (Section 3.2). All these strategies accelerate pre-training by masking specific words in the text.
Consider a text $T_{i}$ consisting of words $\{w_{1},w_{2},w_{3},\ldots,w_{n}\}$, where $n$ represents the total number of words in $T_{i}$. Text masking involves selecting a subset of these words to create a new version of the text, denoted as
$T_{new\_i}=f(T_{i}\{w_{1},w_{2},w_{3},\ldots,w_{n}\}).$ (1)
This function $f$ determines which words are masked to get the modified text $T_{new\_i}$. And, we define the number of input tokens as $k$. It is important to note that the number of padding tokens remains consistent across all these strategies.
@@ 3 Text Masking
3.1 Existing Text Masking Strategies (CLIPA)
The four text masking strategies from CLIPA [12], which we also investigate here, are:
@@ 3.1 Existing Text Masking Strategies (CLIPA)
Random Masking: In the random text masking strategy, the function $f$ operates to randomly select words from the text $T_{i}$. Each word $w_{i}$ has an equal probability of being masked, independently of its position or frequency in the text.
Truncation Masking: In truncation masking, the function $f$ reduces the text by masking words from the end part of the text. The number of text tokens retained depends on the input length of the text encoder.
@@
Block Masking: Block masking is similar to truncation masking but starts at a random index within the text. If the text tokens are longer than the input tokens, we choose a random starting index $j$ that is less than the length of the text tokens minus the length of the input tokens $k$.
@@
Syntax Masking: To implement syntax masking [12], each word $w_{i}$ in the text $T_{i}$ is first tagged using a natural language processing tool like NLTK [23] to identify its part of speech. The words are then sorted according to their syntax role. This ordering strategy $O(w_{i})$ reflects the importance of nouns, adjectives, and verbs in the text. Words are masked based on their assigned order, with priority given to retaining nouns, followed by adjectives, verbs, and then other parts of speech. This method ensures that the core semantic information of the text is less likely to be masked, especially when learning the relationship between the image and the text, as nouns are more relevant to the object of the image.
3.2 Word Frequency Masking (CLIPF)
As mentioned in the introduction, subsampling of frequent words was introduced by [13] in the natural language processing (NLP) domain to accelerate the learning process for word representations. The goals are to speed up training and to address the imbalance between frequent and infrequent words in the training corpus. Our CLIPF method, uses Formula 2, which was introduced in [13], to determine the probability of masking a word during training. Thus, the complete formula to calculate the probability $P(w_{i})$ is:
$P(w_{i})=1-\sqrt{\frac{t}{f(w_{i})}}$ (2)
Here, the masking probability of the word $w_{i}$ is $P(w_{i})$, $t$ is the threshold, and $f(w_{i})$ is the frequency of the word. This formula scales the word frequency $f(w_{i})$, resulting in a much higher masking probability for frequent words than infrequent ones. It also preserves the overall ranking of word frequencies while aggressively masking high frequency words, controlled by the threshold $t$ [13].
4 Experimental Setup
Dataset We pre-train the models on Conceptual Captions 3M (CC3M) [24], Conceptual 12M (CC12M) [15]. These datasets were selected due to their diverse range of real-world concepts and scenes. Popular VLMs such as OpenCLIP, FLIP, DeCLIP, SLIP, A-CLIP and CLIPA [9, 8, 20, 5, 11, 12] also employ these datasets for training, leveraging their extensive and diverse content. Due to the expiration and inaccessibility of some URLs in the training set, we have downloaded 2.7 million image-text pairs for CC3M and 9.3 million image-text pairs from CC12M. The characteristics of the dataset are presented in Table 1. The CC12M dataset is 3.42 times larger than the CC3M dataset, with captions that are on average 2.15 times longer. Since the number of input tokens is fixed, the impact of different text masking ratios varies across datasets due to differences related to caption length. In an additional set of experiments, we provide results training on LAION-400M [17], a larger data set with about 300M image-text pairs.
Dataset Samples Total words Caption length
CC3M 2721877 28022082 10.30 $\pm$ 4.72
CC12M 9295444 205716854 22.15 $\pm$ 17.20
Architecture For the image encoder, we use ViT-B/16 [25] architectures with global average pooling and learnable positional embeddings. The input image size is 224 for CC3M and CC12M. For the text encoder, we use a Transformer-based model [26] and byte-pair encoding with a 49K token vocabulary. We chose these encoders because they are the largest encoders that we can use given our resources. When using full text for training, the maximum context length is 32. Zero-padding is applied to input text that is shorter than the maximum token length of the model.
Table 2: Details of the pre-training and fine-tuning setups.
Config Pre-training Fine-tuning
optimizer AdamW [28] AdamW [28]
learning rate 1e-3 1e-5
weight decay 0.2 0.2
learning rate schedule cosine decay [30] cosine decay [30]
warmup steps 10k 10%
epoch 30 1
$t$ $10^{-6}$
$\tau$ 0.07
numerical precision amp amp
RandomResizedCrop (50, 100) (50, 100)
Image masking Text masking Image tokens Text tokens Total Percentage
0% 0.00% 196 32 228 100%
75% 0.00% 49 32 81 35.53%
50.00% 49 16 65 28.51%
75.00% 49 8 57 25.00%
81.25% 49 6 55 24.12%
87.50% 49 4 53 23.25%
Training Following CLIP, OpenCLIP [3, 9], we pre-train our model for 30 epochs on the CC3M and CC12M datasets. Details of the pre-training configuration are given in Table 2.
Similar to FLIP, to speed up training, we apply 75% image masking to the image encoder as the baseline model. As a result, the speedup is about 4$\times$ compared to training without image masking, while the reduction in the zero-shot performance of ImageNet-1K classification remains within reasonable bounds.
During text masking, we reduce the number of tokens from 32 to 16, 8, 6, and 4. To measure the training speed of CLIP, CLIPA, and CLIPF, we compare the number of text tokens processed by each model. As shown in Table 3, the number of text tokens processed by each model during pre-training is calculated based on different image and text masking ratios. When we pre-trained the model using image masking, the total number of tokens is 81 and the percentage of text tokens is 39.5%. When we apply 50% text masking, the total number of tokens is 65. Compared to training without text masking, this results in a speed increase of approximately 20%. Moreover, when we apply 87.5% text masking, the total number of tokens is 53, resulting in a speed increase of approximately 34% compared to training without text masking.
Fine-tuning Following FLIP, and CLIPA, in order to bridge the distribution gap between pre-training and evaluation, we fine-tuned the model without images and text masking. Note that, fine-tuning the models without masking in masking work is focused on reducing the distribution gap and applied sparingly to avoid undoing the tradeoff. The details of the fine-tuning configuration are provided in Table 2.
Table 4: Comparison of CLIPF, CLIP, and CLIPA for zero-shot classification on ImageNet-1K. Reported as top-1 accuracy. All models use a ViT-B/16 image encoder backbone pre-trained on CC3M and CC12M for 30 epochs with 75% image masking for speedup (pre-train) and fine-tuned for an additional epoch on all data (fine-tune). “Image tokens” and “text tokens” reflect the amount of image and text processing during pre-training.
Models Masking Image Tokens Text Tokens CC3M CC12M
pre-train fine-tune pre-train fine-tune
CLIP 197 32 18.6 36.6
FLIP 49 32 14.1 14.2 32.0 33.7
CLIPA truncation 49 16 13.8 13.8 32.8 32.8
random 13.9 13.9 33.7 34.3
block 13.9 13.9 34.3 34.8
syntax 13.3 12.8 32.2 34.4
CLIPF frequency 14.0 14.0 35.5 36.0
CLIPA truncation 49 8 10.8 12.0 25.4 28.4
random 17.6 17.4 34.5 36.9
block 16.2 16.6 35.5 37.9
syntax 17.2 17.5 28.5 35.0
CLIPF frequency 16.8 17.0 36.6 39.3
CLIPA truncation 49 6 8.4 9.4 15.3 23.2
random 12.8 17.9 26.9 34.6
block 12.9 17.0 28.6 35.9
syntax 12.2 15.7 25.2 32.6
CLIPF frequency 14.4 18.2 30.3 37.8
CLIPA truncation 49 4 3.8 8.2 5.3 19.8
random 5.4 14.6 14.0 27.1
block 7.5 14.5 18.7 26.6
syntax 8.9 13.0 14.2 24.6
CLIPF frequency 10.9 16.0 17.0 30.9
CLIPF frequency 98 8 19.6 18.7 39.8 41.0
5 The Importance of Frequency Distributions
Our work is motivated by the patterns that we observe when VLMs are trained for additional epochs, shown in Fig. 1. This experiment motivates revisiting CLIPA [12], which was trained only on 6.4 epochs, in two respects. First, as already mentioned in the introduction, the conclusion that syntax masking is the best masking approach does not hold. Second, truncation masking, which looks relatively promising in low numbers of epochs, does not perform well given more epochs. In this section, we analyze the impact of masking on the training data in order to discover the reason behind this behavior.
@@ 5 The Importance of Frequency Distributions
We start with a basic illustration of the differences between task-masking strategies is provided in Figure 2. Truncation consistently retains the beginning of the text. With the random (random word selection) and block (random block selection) methods, each word has an equal probability of being masked, with block masking maintaining word neighborhoods. Syntax masking prioritizes the retention of nouns, followed by other part-of-speech (POS) categories.

### v3

# Frequency Is What You Need: Considering Word Frequency When Text Masking Benefits Vision-Language Model Pre-training

@@ Abstract
Vision Language Models (VLMs) can be trained more efficiently if training sets can be reduced in size. Recent work has shown the benefits of masking text during VLM training using a variety of strategies (truncation, random masking, block masking and syntax masking) and has reported syntax masking as the top performer. In this paper, we analyze the impact of different text masking strategies on the word frequency in the training data, and show that this impact is connected to model success. This finding motivates Contrastive Language-Image Pre-training with Word Frequency Masking (CLIPF), our proposed masking approach, which directly leverages word frequency. Extensive experiments demonstrate the advantages of CLIPF over syntax masking and other existing approaches, particularly when the number of input tokens decreases. We show that not only CLIPF, but also other existing masking strategies, outperform syntax masking when enough epochs are used during training, a finding of practical importance for selecting a text masking method for VLM training. Our code is available online.
@@ 1 Introduction
However, pre-training CLIP models typically requires thousands of GPU-days [32, 15]. To accelerate the pre-training process, current work makes use of masking strategies for image and text tokens when pre-training CLIP models [22, 43, 20]. In addition to the speed-up, masking makes it possible to use large batch sizes, which enhances the benefits of contrastive learning and has been considered as a form of regularization, reducing the risk of overfitting [22]. For these reasons, despite the loss of some information from the original image and text, masking strategies have the potential to improve model performance.
@@
Figure 1: Zero-shot classification accuracy on ImageNet-1K using models trained on CC12M to which different text masking strategies have been applied. The backbone of the image encoder is ViT-B/16. The text masking ratio is 75%. We also use image masking (75%) to speed up pre-training. At each point, we apply an additional epoch of training on the full, unmasked data.
Currently, the state-of-the-art in text masking is represented by CLIPA [20]. The authors compare two strategies originally introduced by FLIP [22], truncation and random word selection, and two new approaches, random block selection and syntax-based selection, i.e., masking on the basis of part-of-speech (POS) information. In [20], syntax masking is found to deliver the highest performance.
In this paper, we carry out an analysis of word frequency in the training data before and after text masking and discover a connection between the impact of a masking strategy on the training data and the success of a VLM. As shown in Fig. 1, our experiments reveal that syntax masking falls behind other text masking strategies (random and block) as the number of epochs increases. Our experiments establish two important conclusions: First, syntax masking is not the best performing text masking approach, updating the state-of-the-art finding in [20]. Second, considering word frequency allows us to explain why random and block masking outperform truncation and syntax masking.
This analysis motivates Contrastive Language-Image Pre-training with Word Frequency Masking (CLIPF), a new text masking strategy that directly leverages word frequency proposed in this paper. Because CLIPF leverages word-frequency directly, it does not need to expend compute on POS tagging, necessary for syntax masking.  Fig. 1 shows that CLIPF outperforms syntax masking. CLIPF is inspired by earlier training of text models, that has leveraged frequency masking [29]. The key algorithmic novelty of CLIPF is the use of frequency-based word scores to prioritize words for masking, rather than to impose a hard decision on word removal.
An important aspect of  Fig. 1 is the development of performance over epochs. In contrast to the experiments in FLIP [22] and CLIPA [20], we continue training until additional epochs yield only minimal gains. Fig. 1 demonstrates that our CLIPF approach is useful because it is the top performer in the range of 16-40 epochs, typically used in practice (e.g., OpenCLIP [15] trains on 30 epochs.)
In sum, our paper makes the following contributions:
  • We show that considering the impact of text masking on word frequency in the training data differentiates between successful and less successful text masking strategies.
  • We introduce Contrastive Language-Image Pre-training with Word Frequency Masking (CLIPF), which leverages frequency directly during masking, and perform an in-depth quantitative comparison among existing masking strategies and CLIPF.
@@
Our quantified evaluation among existing alternative text masking approaches is carried out at two scales. A large-scale experiment on the LAION-400M [35] data set reproduces the CLIPA results and shows CLIPF outperforming syntax masking. A large number of smaller scale experiments on the CC3M [36] and CC12M [3] data sets, which are widely used for VLM research [15, 21, 30, 43, 18], demonstrate that consideration of frequency is important in two ways. First, the frequency-based approach CLIPF emerges an excellent choice for text masking, especially as the number of input tokens decreases. Second, the text masking approach should be selected specifically for the application scenario (i.e., data set, input size, number of training epochs). However, in general masking strategies that disproportionately impact word frequency (truncation and syntax) should be avoided and strategies that protect relative word frequencies (random, block and CLIPF) should be preferred. Our code is available online. [^1] https://github.com/MingliangLiang3/CLIPF
@@ 2.1 Vision-Language Models
Vision-language models (VLMs) learn image embeddings from large-scale natural language supervision, which is important for making models transferable. CLIP [32] is the leading example of a VLM widely-used in practice, and other examples are ALIGN, SigLIP, and FLIP [32, 16, 45, 22], CLIP utilizes an image encoder and a text encoder to generate the representation of the image and text, and learn the joint representation by contrastive learning [32, 39]. Because CLIP does not release its pre-training dataset, OpenCLIP pre-trained models on several large-scale public datasets [3, 38, 35, 34] to reproduce CLIP.
Although CLIP and OpenCLIP are pre-trained on very large-scale datasets, recent works like DeCLIP [21] and SLIP [30] demonstrate that Vision-language Models (VLMs) can be trained efficiently with smaller datasets, enabling researchers with limited computational resources to contribute. Moreover, masking image patches or reducing image resolution during VLM pre-training also substantially speed up pre-training while maintaining or enhancing performance, as shown by techniques like FLIP, RECLIP, A-CLIP and CLIPA [22, 19, 43, 20].
@@ 2.2 Text masking for VLM training
Masking text presents another viable option to speed up and enhance the pre-training of VLMs, and is particularly effective alongside strategies that employ a high ratio of image patch masking [22, 20]. FLIP further explored text masking techniques, namely, random and truncation masking (referred to as “prioritized”), to enhance efficiency [22]. While random text masking involves masking text arbitrarily, truncation masking in FLIP strategically eliminates padding tokens and retains the words that preceded them. The study of text masking for VLMs got off to a slow start, since [22] found that text masking promoted efficiency, but deteriorated performance.
@@
Subsequently, CLIPA [20] introduced block masking and syntax masking. As already mentioned in Sec. 1, CLIPA [20] shows that syntax masking achieves the best performance when applying text masking. CLIPA syntax masking prioritizes the retention of nouns, followed by adjectives and then other words. In [20], syntax masking outperforms the truncation, random, and block approaches. However, [20] experiments only with a fixed number of epochs, and does not test performance as the number of training epochs increases. An important contrast with our work, as previously mentioned, is our attention to the number of training epochs. Further, CLIPF, as we will see, counters the danger of giving overly high priority to nouns, leading to overfitting, and also avoids overly suppressing words of the POS category “Other”.
As mentioned in the introduction, CLIPF is inspired by earlier training of text models [29], which selected words by frequency. An initial attempt was made to transfer this idea to CLIP training in a short workshop paper [23] published at the same time as CLIPA. The algorithmic novelty of our CLIPF strategy compared to the frequency masking in  [29] and [23] is the use of the word masking probability score for prioritization, based on the insight that the hard threshold must be dropped. In short, in order to exploit frequency information effectively, each training text needs to fill as many available input tokens as possible, as is done in CLIPF.
3 Considering Word Frequency For Masking
In this section, we introduce existing text masking strategies and carry out an analysis on their impact on word frequency in the training data, which reveals why random and block masking outperform syntax and truncation masking.
3.1 Existing Text Masking Strategies
@@
Here, we define text masking and provide descriptions of the existing text masking strategies introduced by FLIP [22] and CLIPA [20].
Text Masking: Consider a text $T_{i}$ consisting of $n$ words $\{w_{1},w_{2},w_{3},\ldots,w_{n}\}$. Text masking involves selecting a subset of these words to create a new version of the text, denoted as
$T^{\prime}=f(T\{w_{1},w_{2},w_{3},\ldots,w_{m}\}),$ (1)
In the context of VLMs, the result of text masking is a string of tokens:
$T^{\prime}=f(T\{t_{1},t_{2},t_{3},\ldots,t_{k}\}),$ (2)
where $k$ is the length of the input and is determined by the masking ratio. The text is then tokenized, and $k$ tokens are chosen for use in training.
Truncation Masking: In truncation masking, the first $k$ tokens in text $T$ are used as the input.
@@
Random Masking: In random text masking, the function $f$ operates to randomly select tokens from the text $T$. Each token $t_{i}$ has an equal probability of being masked, independently of its position or frequency in the text.
@@
Block Masking: Block masking is similar to truncation masking but starts at a random token index $j$ within the text $T$. If $T$ contains more tokens than there are input tokens, we choose a random starting index $j$ that is less than the number of the text tokens minus the number of the input tokens $k$.
@@
Syntax Masking: To implement syntax masking [20], each word $w_{i}$ in the text $T$ is first tagged using a natural language processing tool like NLTK [2] to identify its part of speech (POS). The words are then sorted according to their syntax role. Priority is given to nouns, followed by adjectives, verbs and then words belong to other POS categories. The first $k$ tokens are used as the input.
An example that illustrates the way that the different forms of masking impact a text in the training set is provided in Fig. 2.
@@
3.2 Impact of masking on word frequency